Posts

Showing posts from April, 2025

NPP Projection Close 4/30

To do this evening/ was completed: T3 intro went smoothly.  Continued downloading NoBudge files in T1 to the NoBudge folder. Just 8, 9 and 10 left. Started transcoding the new file of I Killed My Mother toward the end of the night. Tremors blu-ray in 4/6. Marquee updated with new 35mm letters. Pride and Prejudice down. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 4/30

To do this evening/ was completed: A note for tomorrow: Please see the outgoing section in this blog. Also, friendly reminder to be on the lookout for the BluRay of Cleopatra Jones. I will be at our morning managerial staff trip and have switched shifts with Will, but I'll forward any pertinent info throughout the day.  Weekly update is out! Have a good read when you can, please. Spoons is tomorrow...make sure to prep in the morning, please!  Staff Screenings have been emailed out, playlists built, scheduled on the LMS and blocked on the calendar.  Presentation Issues: Preshow Clips/Playlist: Tech'd the playlist fo Thunderbolts . New comics preshow looks/sounds good. Feature is all set.  Trailer Updates: Special Events/ AV set up: We are all set for the screening of  In the Loop this evening in T3. I tested the mics, playlist, etc. Spots are working, just remember to flip the switch!  Ordering/Receiving (posters & trailers): Ordering/Receiving (content...

NPP Projection Close 4/29

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Double checked and transferred the In the Loop  show playlist to theater 3 and added it to the schedule on Screenwriter   Trailer Updates: Special Events/ AV set up:  The Rules of the Game screening in theater 1 went very smoothly (cleaned the projectors following the show and closed both exhaust ports) Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Downloaded and ingested the "MUBIPreRoll25" pre-show files Kat sent over this evening to the LMS Cristina sent an email this afternoon regarding the delivery of Cleopatra Jones on blu-ray for Shouting at the Screen so please keep an eye out for that if/when it arrives on 5/1 Outgoing/Shipping (content): Re-inspected The Rules of the Game after the show and scheduled a pickup/printed return labels (tracking ID: 880908870209) Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets...

NPP Projection Admin 4/29

To do this evening/ was completed: LMS scheduled for the upcoming week, including Thunderbolts, playlists built, Comic V2 elements ingested. Loops ran for the show tonight. Tech'd Good Night and Good Luck with Desmond, volume logged. Print now on house reels. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Swapped out the lenses in Theater 4 to Flat and focused them Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Tried ingesting the Alice Junior DCP from the physical hard drive Desmond brought over, the ingest still failed unfortunately. My guess is this is an issue with their DCP as I'm sure we got the complete file they uploaded. I let Desmond know. -A Outgoing/Shipping (content): Print/video Inspection: Desmond took a look at Good Night and Good Luck and Postcards From the Edge   Asset tracking doc updated/checked?: Maintenance/cleaning: I'm putting a new XLR connector in the Mic 2 input in Trees, so t...

NPP Projection Report 4/28 Close

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: The Music Driven screening and Q&A went very well tonight. Mics back under the curtain, spots power switched off. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Downloaded the complete Alice Junior DCP but it failed while ingesting. I think there might be an issue with the actual DCP but we can try ingesting it off the drive Desmond brings over tomorrow just to see. Ingested News from Home . Outgoing/Shipping (content): Print/video Inspection: Finished inspecting Postcards from the Edge . It is Dolby A! It looks like mono for a lot of the soundtrack, but I saw some definite distinctions in the waveforms. Overall in decent shape and it's on house reels. Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Filled in all first run on next week's cue sheets. All cinema and private...

NPP Projection Admin 4/28

To do this evening/ was completed: The fire alarm went off this morning whilst I was checking cues in theatre 3. Each theatre was reset and theatre 3 cues are working as normal. Went through all houses for  maintenance checks. - L Finished inspecting Batman and Robin, back on shipping reels for now. Put Rules on house reels for tomorrow's show. Marquee updated with new brunch titles. Reset all lighting panels after the fire alarm went off this morning. Last round in T5 tonight overlaps with the previous. It's on the LMS scheduled 10 min after its actual start time. You can either leave it there and fast forward or do a full manual start when the time comes. I left a note on the 4/6 desk. DHL/Universal contacted for Devil Is A Woman return. Info attached to the cans. This will get picked up tomorrow morning (or, rather, between 10am - 4pm...but likely first thing). - Lillian Robocop ingested to LMS. Presentation Issues: Preshow Clips/Playlist: Playlist set for In the Loop - L Tr...

NPP Projection Close 4/27

To do this evening/ was completed: Inspected the first four reels of Batman and Robin. They're in good condition, back on shipping reels for now. Gave a guest a captiview device for The Shrouds but turns out the DCP does not come with any accessibility features, meaning no HI/VI. So I updated the cue sheets for the rest of the week with N/A for audio/captions. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Admin 4/27

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built and scheduled a special event show playlist for Beware of Mr. Baker tomorrow evening in theater 3 **this still needs credits checked so if cues can be added to it in the morning it would be greatly appreciated + additional set up for the mics and chairs** Filled out the cues for Phenomena in theater 2 tomorrow evening  Trailer Updates: Special Events/ AV set up: Tech'd Bacurau before its screening in theater 3 this morning and removed the feature content from the server and LMS after the show The Devil is a Woman screening in theater 4 went smoothly (cleaned the projectors following the show and turned off the power supplies) Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Re-inspected  The Devil is a Woman and packaged it for return shipping (but did not yet reach out to DHL about its pickup...the canisters have ...

Projection Close 4/26

To do this evening/ was completed: Still no issues in T3.  Dance Girl Dance preshow downloaded and ingested Listening device brought back to its charging port.  Inspected Abominable Dr. Phibes - print report up!  Double checked the A1 playlists.  Cleared content and went through upcoming events for the month of May.  Still waiting on some content for SE playlists.      ` Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content):  Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Report 4/26 Open

I sent an email to Will and Jack about this but just in case: there is a half marathon in Brooklyn tomorrow (the finish line is by us in Prospect Park), please plan ahead to see if this will affect your commute. Carly sent out an email with the road closures. Thank you! To do this evening/ was completed: Did the Wall-E blu-ray get returned to the host on Thursday night? I found the disc (one of three, this is the one that's got the feature on it) in the disc drive down in booth 1. Can't find the case anywhere. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Devil is a Woman went very nicely in T4. Print rewound, projectors cleaned for tomorrow. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): I tried ripping the Tremors blu-ray but it's a 4K disc, we can't work with these with the equipment we have. The disc drive just spits it out. I let Chenelle know. Maybe we will buy a regular blu-ray or download ...

NPP Projection Admin 4/25

To do this evening/ was completed: All sides of the marquee have been updated. Superhuman private screening went smoothly this morning. - A Cues in T3 are functioning normally for now. Showed Jack how to manually adjust the lights on the relay board. If cues stop working again, first try to restart the IMS webpage. - A Alice Jr. DCP files via Dropbox did not download fully, a frequent issue with Dropbox, where it says the download is complete but the file size is in fact incorrect. We're re-downloading the zip file for this DCP, a successful download should be about 94.5 GB. If the download completes and the file size is lower than that then we'll have to figure out another way, or just keep trying until it's successful. Tech'd The Devil is a Woman, The Accountant 2, Until Dawn, and The Rules of the Game. It looks like the P2 1.37 plate will need to be filed slightly as I was catching just a bit of it at the top of the image in T1 when looking at Rules. New forecasts do...

NPP Projection Close 4/24

Image
To do this evening/ was completed: Was kept busy throughout the evening due to the technical issues plaguing theater 3 mentioned in Alec and Lillian's blog needing continuous attention, a private event in theater 7, and a 35mm film screening in theater 1 intro'd by Cristina Presentation Issues: Continued to bring the lights in theater 3 up and down throughout the evening with the paper clip when necessary...hopefully this can be fixed sometime tomorrow but until then... Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up:  Ninja 3: The Domination screening in theater 1 went smoothly and the print was brought downstairs after the show for it to be picked up tomorrow (cleaned the projectors following the show and closed the exhaust ports) WALL-E private event in theater 7 was running smoothly up until the hosts requested the feature be delayed ten minutes since their guests were arriving too slowly...rewinding the show playlist threw the cues off and the film star...

Projection Admin 4/24

Image
To do this evening/ was completed: Had a super fun time in theater 3 today...take a look at Alec's notes. We've both notified Chris/Gregg to get someone in here tomorrow morning to fix the issue.  Friendly reminder that there is a school screening for the Superhuman Body tomorrow morning, right at 10am (film starts around 10:30). Playlist is scheduled but the letters (in 3/7) will need to go up. This will be our last show with those folks! - Lillian  Carefully cleared a lot of content from servers 1, 2, 3 and 4 to prep for our incoming shows and to make transfers a little easier as it was a bit of a struggle before. - Lillian Presentation Issues: We had to move the first round of Sinners from T3 to T1 (and cancel Sacramento ) because the lighting and masking cues in T3 completely stopped working. The masking is easy to get around, you just use the motor under the curtain, but the lighting dimmer requires automation cues to activate the different settings, so we were stuck on ...

NPP Projection Close 4/23

To do this evening/ was completed: Marie Antoinette event went smoothly. All equipment broken down after the show. Pride keys came in, uploaded. T3 shutdown playlist did not strike the lamp off. Wonder if this is related to the recent server issues we've been having with 3. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

Projection Admin 4/23

To do this evening/ was completed: Staff Screenings have been emailed out, blocked on the calendar and LMS. Playlists are all set for tomorrow and Friday.  Presentation Issues: Preshow Clips/Playlist: Superhuman Body playlist all set for Friday on the LMS.  Wall-E playlist for the private event tomorrow is also all set! - Lillian  Trailer Updates: Special Events/ AV set up: Marie Antoinette show is all set for tonight - speakers, mixer and everything all ready to go. - L/ J Downloaded and ingested Having Fun with Assholes to the LMS. Ingested Blues Run the Game to the LMS. FYI - Kat has the DCP of Home Alone 2 to the drive (currently in the doxnotxingest folder) for the private event this weekend. We have the MKV in case something is off with the DCP. - Lillian  Ordering/Receiving (posters & trailers): Ordering/Receiving (content):  *I am downloading the DCP of Alice JR to the iMac in T1 again (network funkiness from last night prevented it from properly ...

NPP Projection Close 4/22

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Added the Marie Antoinette drag show playlist to the schedule on Screenwriter for tomorrow evening Trailer Updates: Special Events/ AV set up: Ran loops between rounds in theater 1 and refocused the lenses for the show tonight +  Waterworld  screening this evening went smoothly Ordering/Receiving (posters & trailers): One sheets for The Shrouds and Final Destination: Bloodlines were delivered in addition to character posters for  One Battle After Another Ordering/Receiving (content): Downloaded and uploaded keys for The Shrouds , Marie Antoinette ,  The Legend of Ochi , and  The Accountant 2 to the LMS Outgoing/Shipping (content): Print/video Inspection: Re-inspected  Waterworld after the show and packaged it for its return/reached out to DHL to pick it up in the coming days (brought down  Babe  too since it is also going to back to NBC) Asset tracking doc updated/...

Projection Admin 4/22

To do this evening/ was completed: Built the cue sheets with new formulas (except Thunderbolts info) for the week of 4/25 - 5/1. - Lillian  LMS scheduled for the coming week with the exception of Thunderbolts. Currently no OCAP Shrouds, so I scheduled the CCAP versions for the listed OCAP screenings, assuming we don't get another DCP. Ran loops for tonight, all set. Tech'd all of the Marie Antoinette content, spots, slides, dance number DCPs and the feature. Volume cues have been added to the playlist. We made a DCP for the curtain call song Primadonna with the breakdown slide attached. Instead of playing the song off of a laptop and then transitioning to breakdown slide in the playlist, we're just going to have the dowser closed during the curtain call bow and once that's concluded I'll open the dowser to display the slide while the song continues to play. Presentation Issues: Preshow Clips/Playlist: FYI - There's no OCAP shows for The Shrouds as there's ...

NPP Projection Close 4/21

To do this evening/ was completed: Re-animator went smoothly. Recorded most of the intro, missed the first minute. Mic back in the booth. Inspected the SoF print, Day Without Policemen. In good condition. Looks a little faded-pink to me but we'll see when we put it on screen. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 4/20

To do this evening/ was completed: Finished re-inspecting La Dolce Vita. Packaged it up for outgoing, scheduled a Ground pickup for tomorrow, print is on the cart by the side entrance. Put up new brunch and 35mm letters. Transferred Re-Animator to T1. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Admin 4/21 (Lillian & Will)

To do this evening/ was completed: Transferred Phenomena to T2 and Beware of Mr. Baker to T3 to take a look at these (they were downloadable links, so I just wanted to double check.) No slide for Baker yet, but the Phemomena playlist has been updated - Lillian  Presentation Issues: Preshow Clips/Playlist: Adjusted all playlists that needed the Dance Girl Dance Series Trailer as opposed to the trailer for the feature as per Kat/Cristina's request this morning. - Lillian  Trailer Updates: Special Events/ AV set up: Tech'd the Re-Animator show playlist this morning and added a house lights up cue for credits -WB Ordering/Receiving (posters & trailers): The Accountant 2 posters were delivered this afternoon and stored behind 3/7 - WB Ordering/Receiving (content): Ingested The Accountant 2 to the LMS, deleted the 5.1. - Lil Outgoing/Shipping (content): Got rid of old playlists/content. La Dolce Vita print outbound. Superbad is getting picked up tomorrow via Sony- L Print/vid...

NPP Projection Report 4/20 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Dolce Vita show was good. Last changeover I started looking at the screen probably half a second too late so mistook the CO cue for the MS cue. Audience saw a bit of tail leader. Pretty peachy other than that. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Reinspected and packed up the first 6 reels of Dolce Vita, the first box. This came with return labels so I attached the first one to the box and taped it up. Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Sent another detailed report from the T3 IMS to Chris from DMS after the web controls went down again yesterday. Cue sheets updated/checked (every shift):  Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 4/19

To do this evening/ was completed: A super quiet night apart from the occasional call to put the cleaning lights back on and putting on the NBA playoffs in Trees.  Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Report 4/19 Open

To do this evening/ was completed: Please put on the Knicks playoffs game in Trees (no sound) tonight at 6! It's ESPN channel 570 (cable box is already set to this channel). I moved the printed reminder to circuit Ninja III to the Trylon Theater from the 4/6 whiteboard and taped it to the box the print is in, in case anyone is wondering where that went. On a related note, printed the email from Chris at Park Circus to return To Live and Die in L.A. to P3 immediately after its show (which is not until late July). This is posted on the wall above the T1 rewind bench. Presentation Issues: Theater 3 server 404'd on me today. I power cycled the projector and got the preshow on about 15 minutes late (skipping ahead so the feature started on time, of course). This was before the 3pm Sinners . Keep an eye on it, though it shouldn't have any more issues tonight. Preshow Clips/Playlist: I deleted the old happy hour PSA (Steve_V1) from all servers as we should only be using the one ...

NPP Projection Close 4/18

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Received one sheets for The Naked Gun  and an additional teaser poster for Mission: Impossible - The Final Reckoning  and stored them behind 3/7 Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Spent the majority of the evening finishing inspecting the  Barber Shop print in booth 1 and rewound it so all reels are tails out on their shipping reels Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Yes Additional Notes:

Projection Admin 4/18

To do this evening/ was completed: Tech'd Children of Men  and updated info in the playlist and on the cue sheets. I also added in the credits info for Babe to the cue sheets. All set for this weekend. - Lillian  Tech'd Waterworld, volume logged. Re-inspected Superbad, all packaged up, can reach out to Sony soon to coordinate pickup.  Brought La Dolce Vita down to 1. All reels are heads out. Started inspecting Barbershop. First two reels are done. Presentation Issues: Preshow Clips/Playlist: All new A1 playlists have been built for the week of 4/25, including every SE show apart from Beware of Mr. Baker (waiting for event memo) and In the Loop (waiting for slide) - Lillian  Trailer Updates: Ingested new trailers and built all new trailer packs for this next week. - Lillian  Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Received the new content from Kat. Downloaded and ingested the new forecasts. We alrea...

NPP Projection Report 4/17 Mid (Alec) & Close (Will)

To do this evening/ was completed: Presentation Issues:  Preshow Clips/Playlist: Built Re-Animator and Marie Antoinette playlists (A) Trailer Updates: Special Events/ AV set up: NewFest event went very well. Cleared the feature off the servers after it was done. All projection related equipment has been put away. (A) La Dolce Vita screening in theater 4 this evening went very smoothly (save for a moment at the top of reel 5 where the image was out of frame, despite my being confident I'd framed the countdown correctly which required a quick correction) despite it being a busy show with lots of reel changes  **I have left these reels tails out in 4/6 and will rewind them tomorrow in booth 1 since they will be screened there on 4/20** Ordering/Receiving (posters & trailers): Teaser one sheets for HIM were delivered this afternoon and stored behind 3/7 -WB Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Finished inspecting Waterworld , it'...

Projection Admin

To do this evening/ was completed: Loops ran for Dolce. Projectors cleaned afterward. Inspected the first two reels of Waterworld. Nice condition. On house reels. New 35mm letters up, Sinners up too. Presentation Issues: Preshow Clips/Playlist: Tech'd Pride & Prejudice and Sinners . Went over the credits info with Gil and decided on 6 minutes, so I adjusted the playlists and the cue sheets. - Lillian  Trailer Updates: Special Events/ AV set up: NewFest is all set for this evening. Double spots for the Q&A, single left spot (which I programmed again this morning just to position it a little better) has been added to the spots sheet in T1. Sound for preshow and exit is set - iMac has the spotify playlist up and sound has been tested. Should be a super smooth evening. Lillian Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Reached out to Deluxe to expedite new keys for Sacramento  to us as they'll expire tomorrow. We received these. Also, just ...

NPP Projection Report 4/16 Close

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: PnP, Shrouds, and Ochi playlists built and scheduled. Scheduled PnP staff screenings as well Some of the schedules for next week were for the wrong times, be careful of that 24hr clock! Made another version of La Dolce Vita for the weekend, scheduled it. Trailer Updates: Special Events/ AV set up: Superbad went nicely in T1. Rewound the print for the staff show on Friday. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested Pride and Prejudice (the partial OCAP file has been ingested and was all that was showing up on the FilmVault, so I re-scanned the FilmVault and ingested the full CCAP version) Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Last round in T6 had The Friend on the cue sheet but it was supposed to be Death of a Unicorn . The correct playlist was sc...

Projection Admin 4/16

To do this evening/ was completed: Hey y'all - Turns out Pride and Prejudice wasn't booked until today, but we were able to get the feature content and keys in a couple hours! Keys open at midnight, so we're all set for the staff screening tomorrow/Friday. I'll take a look at both Sinners/P&P tomorrow.  *Weekly update is OUT please get your eyes on it when you can! Thanks! - Lillian  Inspected the last reel of Good Night and Good Luck. Ran loops for tonight's show, projectors cleaned afterward. Presentation Issues: Preshow Clips/Playlist: Sinners staff screening playlist all set. All staff screenings have been blocked on the calendar and emailed out, but we need to make a playlist for Pride and Prejudice (it's still ingesting) - Lillian Trailer Updates: Special Events/ AV set up: Built the playlist and ran through all media for the NewFest screening of The Wedding Banquet tomorrow in T1. Spots/Spotify playlist is all set. Content looks and sounds great. Ev...

NPP Projection Close 4/15

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Feature content for  The Shrouds , Pride & Prejudice , and Re-Animator  has not yet been made available so these playlists will have to wait to be built  No pre-show content information for The Legend of Ochi  has been added to the trailer lists but I have reached out to Kat to ask about whether or not an all ages pre-show can be used in place of custom content Built and scheduled the pre-show playlist for Babe  on 4/19 (cues can be double checked later on + the trailer for The Emperor's New Groove  should be tech'd for volume soon too)   Trailer Updates: Special Events/ AV set up: Set up for trivia at Trees Lounge Miracle Mile screening in theater 3 this evening went smoothly Ordering/Receiving (posters & trailers): Fight or Flight posters were delivered and stored behind 3/7 Ordering/Receiving (content): Downloaded and ingested the remaining "Nostalgia_Ultra_V2" c...

NPP Projection Admin 4/15

To do this evening/ was completed: Presentation Issues:  Preshow Clips/Playlist: Playlists for Legend of Ochi, The Shrouds , Pride and Prejudice , Marie Antoinette , Re-Animator , and Babe still need to be built I started making DCPs of the slides for Marie Antoinette . I labeled them SLIDE_MarieAntoinette_[Order #]_Performer/Info to try to make it as uniform and easy as possible. I made the intro slide, Malice, and Mauve's slides. Breakdown and curtain call still need to be done and we're waiting on  Trailer Updates: Special Events/ AV set up: Not so much an NPP thing, but continuing some NWB work, I started the edit of Eras for the private event there Friday. Made the cuts, now I just gotta cobble it all together, smooth out the transitions, and export it.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Started ingesting The Legend of Ochi onto the LMS Keys for Wedding Banquet , Death of a Unicorn , and Sinners uploaded Ingested Marie ...

NPP Projection Close 4/14

To do this evening/ was completed: Save for the event in theater 2 this evening it was a slower night overall, with a handful of manual starts/negative flips requiring lights to be brought up and shows to be played remotely Presentation Issues: Preshow Clips/Playlist: Added a feature volume cue to the Miracle Mile show playlist but this still needs to be tech'd for a lights cue tomorrow morning Trailer Updates: Special Events/ AV set up: Sunset Boulevard  pre-show Q&A and screening in theater 2 this evening went smoothly (mics were returned to the booth to charge and chairs were placed back in the hallway following the introduction) Fright Night played to a smaller crowd in theater 1 without any issues Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Downloaded and ingested "PS_SE_Nostalgia_Ultra_V2" content (25min and 15min, the 5min folder had file elements missing) to the LMS/theater 1 for the Babe pre-show on 4/19 Outgoing/Shipping (conte...

Projection Admin 4/14

Image
To do this evening/ was completed: Forgot to post my blog last night, but no issues. Relatively quiet night on my end. AC units are all fine on the ground floor. Also, the print of Good Night and Good Luck has been all sorted. Ready to be put on house reels when the time comes. Additionally, photographs have been uploaded to the drive for both the aforementioned print as well as Dance, Girl, Dance . Presentation Issues: Preshow Clips/Playlist: Just a friendly reminder to make our SE slides about 20 seconds long. Allows us to add a nice fade in/out, looks nicer on screen, more well paced (gives the other cues more space). Thanks! Here's an example using the Babe slide. It's 20 seconds long (reaches maximum brightness at 8 (maybe between 8 - 10) seconds in, so I like to put the server pause cue there. Fade ins/outs are added in at 5 seconds on each end. See images below. Make sure you click "set" on each Trailer Updates: Special Events/ AV set up: Tonight we have the...

NPP Projection Admin 4/13

To do this evening/ was completed: Presentation Issues: There was some minor confusion with the  God of Gamblers Returns  soundtrack during its first reel. I began running the print with the processor set to Dolby SR, as noted on the inspection report, until Cristina reached out to mention it sounded like it was wavering. I switched to Dolby A, checked in on the auditorium, and left the rest of the reels running in this format since nothing sounded out of the ordinary to me and Cristina said it sounded normal too Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Tech'd the Hyenas show playlist in theater 4 before the first round and adjusted the 1.66 lens file to sharpen focus on the film's subtitles Ran loops in booth 1 before the Sundays on Fire event  God of Gamblers Returns intro and screening went smoothly despite the soundtrack issue (brought the mic back up to the booth to charge and cleaned both projectors after the show + shut the 35mm exhaust v...

Projection Close 4/12

To do this evening/ was completed: Lots of print inspection. Quiet night overall. Presentation Issues: Preshow Clips/Playlist: Checked the new playlists for this upcoming week. Double checked this week's SE shows.  Trailer Updates: Special Events/ AV set up: Both DCPs from Agfa finished downloading/unzipping. Ingested both of them: Phenomena and Re-animator to the LMS.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Received the keys for The Wedding Banquet at midnight!  Outgoing/Shipping (content): Print/video Inspection: Inspected all of Dance, Girl, Dance . Print is in good condition but has a couple of rough-ish patches, but it'll still look lovely. Report is up.  Almost finished inspected all of Good Night and Good Luck . I have one reel left. Print is basically in immaculate condition. Report almost complete. Asset tracking doc updated/checked?: Maintenance/cleaning: Our very cute alcohol dispensers were delivered along with some po...

NPP Projection Report 4/12 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Spoons setup was smooth, broke everything down and returned the mics and spotlight board to the booth afterward. Gosford Park tech'd and SPL cue'd before the show. Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Downloaded the Re-Animator DCP and unzipped it. Phenomena is downloading now. These are on the "George Ivan" hard drive and can be ingested to the LMS directly from that. Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Added some May 35mm titles to the whiteboard since we already have two of those prints in. Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 4/11

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Built and scheduled show playlists for Sinners  on Thursday 4/18 Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Uploaded keys for Sunset Boulevard ,  Gosford Park , and  Fright Night Outgoing/Shipping (content): Scheduled a return pickup for The Original Kings of Comedy  for Monday 4/14 and brought the canisters with labels attached downstairs Print/video Inspection: Finished inspecting Superbad  and placed it on house reels + took a bunch of stills and added those to the shared drive Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Double checked all cues for the weekend, cross referenced the schedule on Screenwriter, and initialed for confirmation Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by entrance?: Yes Addit...

NPP Projection Report 4/11 Admin

To do this evening/ was completed: Please remember to close the exhaust ducts for the 35mm projectors in T1 after shows. Give the lamphouses 15 minutes to cool down then close the duct, this is required as it significantly weakens the airflow from the digital projector lamphouse when these are left open. They were still open this morning. I sent the video I took of the equipment under the curtain in T1 again, please watch it. Presentation Issues: Preshow Clips/Playlist: I switched all the R rated playlists on the 17th to the 4/18 versions ( Sinners opening early, and those 4/11 playlists have the trailer) Built the playlist for Sunset Blvd . Built 4/18 first run playlists with the exception of Sinners and Pride & Prejudice . Sinners can get done tonight. Trailer Updates: Special Events/ AV set up: Went ahead and put Babe on house reels. -If you're inspecting a print but leaving it on shipping reels, it's nice for your fellow projection people to wind it tails out on the...

NPP Projection Close 4/10

To do this evening/ was completed: Presentation Issues: Spoons got off to a delayed start due to a missing power strip (normally placed beneath the curtain where the cereal bar/lights cord is set up for guests) that needed to be replaced and some confusion regarding an XLR cable connecting to the light mixer in the auditorium that was not connected in the booth Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Went over the new HDMI setup in booth 1 with Alec prior to the STB event this evening Spoons ran smoothly despite the aforementioned delay and all gear was placed back beneath the curtain for Saturday morning's brunch show ( note: I placed a tablecloth over the table with the light mixer on it and moved it to the right of the screen and set the mic mixer back beneath the curtain) Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Finished inspecting the remainder of La Dolce Vita an...
To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist:Quick FYI - spent some time hunting down KDMs for SE shows today, and, the KDM for Gosford Park will come in tomorrow, but doesn't open until the morning of 4/12 (which is the day it screens). It should be active starting at 8am, so on Saturday night, I'll leave a note in the blog for the Sunday opener to tech this first thing in the morning! - Lillian  Warfare and Sacramento playlists/preshow and trailers have been techd/updated. - Lillian Trailer Updates: Special Events/ AV set up: Setup for Spoons tonight. Mics/laptop image and sound have all been tested. The Aja, hdmi adapter, and spotlight board are all organized on the table next to the curtain, ready for use. Please just remember to (use Digital 4 on the processor!) and turn on the breakers for the spots before the show along with the front panels of the wireless mic receiver and power switch on the yamaha mixer.  - Lillian  Lori from...

NPP Projection Close 4/9

To do this evening/ was completed: Inspected reels 2 - 4 of Dolce, on house reels. OG Kings of Comedy went smoothly. Print heads out for now. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Admin 4/9

To do this evening/ was completed: Staff Screenings have been emailed out, scheduled on the LMS and blocked on the calendar. Weekly projection update is also out, so take a good look at that when you're available! There's some important housekeeping notes concerning house maintenance, scheduling, expectations, etc. - Lillian   We had our cinema meeting and everything's looking good, we just need Superbad  (coming via Sony) and the NewFest content (coming via Monisha and her team, hopefully as soon as possible as there's a chance we might be DCPing their assets and feature content). We will keep you updated. We're in conversation with them. - Lillian  Presentation Issues: Preshow Clips/Playlist: Built the rest of the SE playlists for this week, apart from Sunset Boulevard , which we'll be going through this Friday. Sacramento playlists sorted. Finished scheduling them on the LMS (just not the overlapping shows) - Lillian  Trailer Updates: Special Events/ AV set...

NPP Projection Close 4/8

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Damsels in Distress show went smoothly. Brought the mic back up to the booth to charge. Cleaned both projectors and switched the ventilation back to "digital" after the lamps cooled Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Re-inspected  Damsels in Distress following the screening, scheduled a return pickup for tomorrow morning (tracking ID: 880407591301), and brought the print downstairs Print/video Inspection: Began inspecting La Dolce Vita , started an inspection report spreadsheet, and added some stills to the shared drive Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Double checked times for tomorrow, cross referenced the schedule, and initialed for confirmation Marquee updated?: Posters updated?: Cleaning lights on for the night?: Ye...

NPP Projection Admin 4/8

To do this evening/ was completed: LMS scheduled for the coming week. Sacramento and Children of Men ingested. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Went through the playlist for Damsels.  Spoke to Chenelle (she was in house today!)about three upcoming private events so we could make a game plan for each - Superhuman Body, NewFest, Live Streaming Event for September (we've got a while). - Lillian  We received multiple DCPs today - links for Bacurau (which I am downloading right now with filemail app and the other, Beware of Mr. Baker , can get downloaded after Bacurau . - Lillian Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested Children of Men . Ingested Hyenas to T4 and tech'd this - cue sheets have been updated. DCP of Hyenas is gorgeous! Reached out to Deluxe, and then Matthew and Jeffrey about Sacramento as it looked like our booking hadn't gone through, but it's all sorted now. -...

Projection Close 4/7

To do this evening/ was completed: Quiet night on our end, all in all.  Received the film schedule a couple minutes ago, so I will lock in with that tomorrow morning (I'm on admin tomorrow, Alec is at WB!) Presentation Issues: Preshow Clips/Playlist: Double checked the playlist for Return of the Living Dead, which played to a happy crowd tonight.  Got rid of old playlists and old content from over the weekend.  Trailer Updates: Special Events/ AV set up: Took a closed captioning device out for T5. Back charging.  Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested Miracle Mile to the LMS.  Ingested Sunset Boulevard to the LMS.  Outgoing/Shipping (content):  Print/video Inspection: SOF print came in, brought it down to T1. Inspected the print. In good condition despite some edge damage (particularly in reel 4). Color is great. Report is available, I have a few pics, but since it's SOF, I'll keep em until requested.  ...

NPP Projection Admin 4/7

To do this evening/ was completed: Rec-scheduled an express pickup today for Westward the Women. Was picked up this afternoon. Looked at Return of the Living Dead. Lights up and volume cues set. 6.0 sounds good to me. We received our print of Babe, which I inspected in full. Good condition. Back on shipping reels for now. One of the shipping reels came damaged, just noting that here. SoF print also arrived. Uploaded the NSF NoBudge and Closing Nite Q&As to Google Drive. Cleaned port-glass windows in 1. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 4/6

To do this evening/ was completed: Presentation Issues: The round of Minecraft  preceding the shorts program left the servers with a pretty substantial mess to clean up after, so the MODs asked that I delay the pre-show to the latter by fifteen minutes, which in turn caused us to have to push back the scheduled showtime for the final round in this auditorium after the NSF filmmaker Q&A ended. Thankfully there were very few tickets sold to Black Bag  and the show played normally, albeit with a fifteen minute delay, after it was started manually. Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: NSF Closing Nite shorts program went smoothly (despite the aforementioned delay to the pre-show) and the Q&A was recorded and saved to the rack desktop in booth 1 + mics from the Q&A were returned to the booth to charge Set up a mic at Trees for the awards announcements and helped John prepare the turntables for Ariel's afterparty DJ set (returned all gear t...

NPP Projection Open 4/6

To do this evening/ was completed: NSF NoBudge went smoothly. Q&A recorded, will transfer to Drive soon. Downloaded and ingested the Sunset BLVD slide as well as a short trailer for Sunset. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Marquee updated?: Posters updated?: Cleaning lights on for the night?: Projection cart by entrance?: Additional Notes:

NPP Projection Close 4/5

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Minecraft  played just fine at 6.0 this evening (fortunately the crowds did not seem too rowdy or eager to make noise of their own over the film) but I will continue to keep an ear on levels Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Westward the Women is still on the cart awaiting its return pickup Print/video Inspection: Finished inspecting The Rules of the Game and added more stills to the shared drive Asset tracking doc updated/checked?: Maintenance/cleaning: Regular garbage and recycling were tossed + placed empty/additional bags in the bins in 4/6 Cue sheets updated/checked (every shift): Double checked all cues and cross referenced the LMS schedule, adjusted accordingly, and initialed for confirmation Marquee updated?: Posters updated?: Cleaning lights on for the night?: Yes Projection cart by...

NPP Projection Report 4/5 Open

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: With how full these Minecraft screenings are getting, I put a 6.0 volume cue for the feature in all playlists. Could even go higher, depending on the crowd. Deleted all 3/28 playlists Trailer Updates: Special Events/ AV set up: NSF Matinee Two went well! Q&A was recorded and it's in the Google Drive folder here Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Ingested the 5.1 OCAP version of Warfare . I don't think we should play the IAB version, not sure if it would sound bad, it's similar to Atmos. There's no regular 7.1 OCAP version. I deleted the IAB one and put the 5.1 version in the OCAP playlist. Outgoing/Shipping (content): Deleted Eephus , Opus , and Flow from everywhere. Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Swept and mopped the booth 2 floor. Cue sheets updated/checked (every shift): Marquee updated?: Pos...